OpenStudy (anonymous):

A line passes through the point (0, 2) and has a slope of -1/2. What is the equation of the line? A. 2x + y = 4 B. x – 2y = 2 C. x – 2y = 4 D. x + 2y = 4

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] y-2=(-1/2)(x-0) y-2=-x/2 2(y-2)=-x 2y-4=-x x+2y-4=0 x+2y=4 its D plz dont forget to click on best response